### v2

@@ 1.1 Inverse problems in the space of Radon measures ${\mathcal{M}(\Omega)}$
We focus now on the formulation of linear inverse problems in ${\mathcal{M}(\Omega)}$. As an acquisition space we will consider a Hilbert space $\mathcal{H}$. Let $\mu\in{\mathcal{M}(\Omega)}$ be the unknown source measure. We consider an acquisition $\bar{y}\in\mathcal{H}$ being the result of the action of the forward operator $\Phi:{\mathcal{M}(\Omega)}\rightarrow\mathcal{H}$ evaluated on $\mu$, i.e. $\bar{y}=\Phi\mu$. The forward operator $\Phi:{\mathcal{M}(\Omega)}\rightarrow\mathcal{H}$ is defined in terms of a a continuous and bounded measurement kernel $\varphi:\Omega\rightarrow\mathcal{H}$, that is:
@@
In the following, we endow ${\mathcal{M}(\Omega)}$ with the weak* topology, making $\mathscr{C}(\Omega,\mathbb{R})$ its dual space. Moreover, we assume that the forward operator $\Phi$ is weak* continuous. Thus, it is possible to define the adjoint operator of $\Phi$ in the weak* topology, namely the map $\Phi^{*}:\mathcal{H}\longrightarrow\mathscr{C}(\Omega,\mathbb{R})$ such that, for all $p\in\mathcal{H}$, $\Phi^{*}p$ is the real continuous function defined by
@@ 2 Off-the-grid Poisson inverse problems
means that for each $m=1,\ldots,M$, the $m-$th element $y_{m}$ is a random realisation of a uni-variate Poisson random variable with mean $(Ax)_{m}+\epsilon_{m}$ depending, in particular, on the unknown vector $x$ [^1] Note that the extension from a discrete support to the real-line can be done using the Gamma function.. Moving from a discrete to a continuos settings requires some attention, since there is no clear infinite-dimensional interpretation of (18).
@@ 2.2 Dual problem
The study of dual problems of sparse-regularisation models with non-negativity constraints has been carried out in [40, 41] in the discrete setting of LASSO, and in [42] for the discrete counterpart of ($\tilde{\mathcal{D}}_{KL}-|\cdot|$), that is with the Kullback-Leibler divergence as fidelity and the $L^{1}$ penalty. In the following we compute the dual problem of ($\tilde{\mathcal{D}}_{KL}-|\cdot|$) in ${\mathcal{M}(\Omega)}$, endowed with the weak* topology by means of $F^{*}$ (32) and $G^{*}$ (31), exploiting the Fenchel-Rockafellar duality. For readability purpose, we report here its statement with coherent notation.
@@
(Fenchel-Rockafellar duality.) Let $V,Y$ be two Banach spaces. For $\Lambda:V\rightarrow Y$ linear and continuous operator, $F:V\rightarrow\mathbb{R}$ and $G:Y\rightarrow\mathbb{R}$ convex functionals, we consider the following primal problem:
@@ 2.3 Extremality conditions
In the extremality conditions (50), the subdifferential $\partial\big{(}|\mu_{\lambda}|(\Omega)+\iota_{{\mathcal{M}^{+}(\Omega)}}(\mu_{\lambda})\big{)}$ requires attention. Recall that, in general, for two proper convex functions $f_{1},f_{2}:\mathcal{X}\rightarrow\mathbb{R}\cup\{+\infty\}$ there holds $\partial(f_{1}+f_{2})\subseteq\partial f_{1}+\partial f_{2}$, with the equality holding if, for instance, $\text{dom}(f_{1})\cap\text{int}(\text{dom}(f_{2}))\not=\emptyset$, see, e.g., [43, Corollary 16.38 (ii)]. In our case, it is possible to show that
$\partial F(\cdot)\subset\partial|\cdot|(\Omega)+\partial\iota_{{\mathcal{M}^{+}(\Omega)}}(\cdot)$ (53)
is a strict inclusion, and to directly compute the analytical expression of $\partial F$.
@@
Given the penalty term $F$ defined in (24), its subdifferential can be directly computed for all $\mu\in{\mathcal{M}(\Omega)}$ and it reads as
@@
$\langle\psi,\bar{\mu}-\mu\rangle=\langle\psi,\bar{\mu}\rangle-\langle\psi,\mu\rangle\leq\langle\psi,\bar{\mu}\rangle-\langle\psi,\mu\rangle.$ (62)
@@ 5.1 Simulated 1D experiments
Using the same dataset, we then compare the results obtained running the homotopy algorithm (Alg.2) for the automatic selection of the regularisation parameter $\lambda$ for both problems ($L^{2}-|\cdot|$) and ($\tilde{\mathcal{D}}_{KL}-|\cdot|$) solved by Algorithm 1 as inner solver, with the algorithmic parameters specified in Table 2. To reduce the computational burdens, we observed that only one iteration of SFW was enough, as the estimated measures are anyway updated in the next homotopy step. For the same reason, we set the maximum number of homotopy outer iterations to be equal to twice the number of peaks in the ground truth. As far as $\sigma_{\text{target}}$ is concerned, being in a simulated environment, we computed the exact value of the residual $\sigma_{\text{exact}}=f_{y^{\delta},b}(\Phi\mu_{\text{gt}})$ and set $\sigma_{\text{target}}=1.5\cdot\sigma_{\text{exact}}$, to be compared with the residual $\sigma_{t}$ at current iteration $t$ of Algorithm 2. Since $\sigma_{\text{exact}}$ is unknown in real situations, a possible strategy for its estimation will be discussed in the next section.
@@
In Table 3, we report the values of TP, FN, FP, Jaccard index and RMSE of the reconstructed signals for both models. The final estimated $\lambda$ is also reported in Table 3, together with the number of performed homotopy iterations and of the value $\sigma_{\text{target}}$. Using homotopy, we observe that we retrieve values which are comparable with the best ones obtained using SFW with grid search. This shows the effectiveness of the homotopy strategy. Overall, the algorithm applied to solve ($\tilde{\mathcal{D}}_{KL}-|\cdot|$) yields better results than ($L^{2}-|\cdot|$) in the presence of Poisson noise, with a reduction of the number of FN and an improvement of the accuracy in terms of Jaccard index.
@@ 5.2 Simulated 2D and 3D examples: choice of $\sigma_{\text{target}}$
Table 4: Different estimates of $\sigma_{\text{target}}$ in the 2D simulated setting.

### v3

@@ 1.1 Inverse problems in the space of Radon measures ${\mathcal{M}(\Omega)}$
We focus now on the formulation of linear inverse problems in ${\mathcal{M}(\Omega)}$. As an acquisition space we will consider a functional Hilbert space $\mathcal{H}$. Let $\mu\in{\mathcal{M}(\Omega)}$ be the unknown source measure. We consider an acquisition $\bar{y}\in\mathcal{H}$ being the result of the action of the forward operator $\Phi:{\mathcal{M}(\Omega)}\rightarrow\mathcal{H}$ evaluated on $\mu$, i.e. $\bar{y}=\Phi\mu$. The forward operator $\Phi:{\mathcal{M}(\Omega)}\rightarrow\mathcal{H}$ is defined in terms of a a continuous and bounded measurement kernel $\varphi:\Omega\rightarrow\mathcal{H}$, that is:
@@
In the following, we consider ${\mathcal{M}(\Omega)}$ with the weak* topology and its dual space $\mathscr{C}(\Omega,\mathbb{R})$. In this setting, the forward operator $\Phi$ is weak*-weak continuous [12]. Thus, it is possible to define the adjoint operator of $\Phi$ in the weak* topology, namely the map $\Phi^{*}:\mathcal{H}\longrightarrow\mathscr{C}(\Omega,\mathbb{R})$ such that, for all $p\in\mathcal{H}$, $\Phi^{*}p$ is the real continuous function defined by
@@ 2 Off-the-grid Poisson inverse problems
means that for each $m=1,\ldots,M$, the $m-$th element $y_{m}$ is a random realisation of a uni-variate Poisson random variable with mean $(Ax)_{m}+\epsilon_{m}$ depending, in particular, on the unknown vector $x$ [^1] Note that the extension from a discrete support to the real-line can be done using the Gamma function.. Moving from a discrete to a continuous settings requires some attention, since there is no clear infinite-dimensional interpretation of (18).
@@ 2.2 Dual problem
The study of dual problems of sparse-regularisation models with non-negativity constraints has been carried out in [40, 41] in the discrete setting of LASSO, and in [42] for the discrete counterpart of ($\tilde{\mathcal{D}}_{KL}-|\cdot|$), that is with the Kullback-Leibler divergence as fidelity and the $L^{1}$ penalty. In the following we compute the dual problem of ($\tilde{\mathcal{D}}_{KL}-|\cdot|$) in ${\mathcal{M}(\Omega)}$, endowed with the weak* topology, by means of $F^{*}$ (32) and $G^{*}$ (31), exploiting results on duality that can be found in [43] and that are similarly applied in the context of BLASSO in [3, 19]. For readability purpose, we report the statement of the needed result with coherent notation.
@@
Let $V$ be a locally convex vector space and let $Y$ be a Banach space. For $\Lambda:V\rightarrow Y$ linear and continuous operator, $F:V\rightarrow\mathbb{R}$ and $G:Y\rightarrow\mathbb{R}$ convex functionals, we consider the following primal problem:
@@ 2.3 Extremality conditions
To have a complete analytical expression of the extremality conditions (50), we compute in the following proposition the subdifferential $\partial\big{(}|\mu|(\Omega)+\iota_{{\mathcal{M}^{+}(\Omega)}}(\mu)\big{)}$.
@@
The subdifferential of the penalty term $F$ in (24) can be directly computed for all $\mu\in{\mathcal{M}(\Omega)}$ as
@@
$\langle\psi,\bar{\mu}-\mu\rangle=\langle\psi,\bar{\mu}\rangle-\langle\psi,\mu\rangle\leq{\color[rgb]{1,0,0}\langle 1,\bar{\mu}\rangle-\langle 1,\mu\rangle}.$ (61)
@@
Remark 2.
Recall that, in general, for two proper convex functions $f_{1},f_{2}:\mathcal{X}\rightarrow\mathbb{R}\cup\{+\infty\}$ there holds $\partial(f_{1}+f_{2})\subseteq\partial f_{1}+\partial f_{2}$. Note that by direct computation there holds
$\partial F(\cdot)=\partial|\cdot|(\Omega)+\partial\iota_{{\mathcal{M}^{+}(\Omega)}}(\cdot).$ (62)
Indeed, the subdifferential of the TV norm is
$\partial|\mu|(\Omega)=\{\psi\in\mathscr{C}(\Omega,\mathbb{R})|\|\psi\|_{\infty}\leq 1\ \text{and}\ \int_{\Omega}\psi\mathrm{d}\mu=|\mu|(\Omega)\}$ (63)
and, the subdifferential of $\iota_{{\mathcal{M}^{+}(\Omega)}}$ can be computed similarly as in Proposition 2 and is equal to
$\partial\iota_{{\mathcal{M}^{+}(\Omega)}}(\mu)=\begin{cases}\emptyset&\mu\notin{\mathcal{M}^{+}(\Omega)}\\ \{\psi\in\mathscr{C}(\Omega,\mathbb{R})|\psi(x)\leq 0\ \forall x\in\Omega\ \text{and}\ \psi(x)=0\ \forall x\in\text{supp}(\mu)\}&\mu\in{\mathcal{M}^{+}(\Omega)}\end{cases}.$ (64)
If $\mu\notin{\mathcal{M}^{+}(\Omega)}$, we have $\partial\iota_{{\mathcal{M}^{+}(\Omega)}}(\mu)=\emptyset$ and $\partial F(\mu)=\emptyset$. Hence, $\partial|\mu|(\Omega)+\partial\iota_{{\mathcal{M}^{+}(\Omega)}}(\mu)=\partial|\mu|(\Omega)+\emptyset=\emptyset$, which is equal to $\partial\big{(}|\mu|(\Omega)+\iota_{{\mathcal{M}^{+}(\Omega)}}(\mu)\big{)}=\partial F(\mu)=\emptyset$.
If $\mu\in{\mathcal{M}^{+}(\Omega)}$, having $\int_{\Omega}\psi\mathrm{d}\mu=|\mu|(\Omega)$ for $\psi\in\mathscr{C}(\Omega,\mathbb{R})$ is equivalent to requiring $\psi(x)=1$ for all $x\in\text{supp}(\mu)$ since $|\mu|(\Omega)=\langle 1,\mu\rangle=\int_{\Omega}1\mathrm{d}\mu$. Thus, the expression of the subdifferential $\partial|\mu|(\Omega)$ reduces to
$\partial|\mu|(\Omega)=\{\psi\in\mathscr{C}(\Omega,\mathbb{R})|-1\leq\psi(x)\leq 1\ \forall x\in\Omega\ \text{and}\ \psi(x)=1\ \forall x\in\text{supp}(\mu)\}.$
Thanks to the above expression, it is straightforward to show that $\partial F(\mu)=\partial\big{(}|\mu|(\Omega)+\iota_{{\mathcal{M}^{+}(\Omega)}}(\mu)\big{)}=\partial|\mu|(\Omega)+\partial\iota_{{\mathcal{M}^{+}(\Omega)}}(\mu)$ for all $\mu\in{\mathcal{M}(\Omega)}$.
@@ 5.1 Simulated 1D experiments
Using the same dataset, we then compare the results obtained running the homotopy algorithm (Alg.2) for the automatic selection of the regularisation parameter $\lambda$ for both problems ($L^{2}-|\cdot|$) and ($\tilde{\mathcal{D}}_{KL}-|\cdot|$) solved by Algorithm 1 as inner solver, with the algorithmic parameters specified in Table 2. To reduce the computational burdens, we observed that only one iteration of SFW was enough, as the estimated measures are anyway updated in the next homotopy step. For the same reason, we set the maximum number of homotopy outer iterations to be equal to twice the number of peaks in the ground truth. As far as $\sigma_{\text{target}}$ is concerned, being in a simulated environment, we computed the exact value of the residual $\sigma_{\text{exact}}=f_{y^{\delta},b}(\Phi\mu_{\text{gt}})$ and set $\sigma_{\text{target}}=1.5\cdot\sigma_{\text{exact}}$, to be compared with the residual $\sigma_{t}$ at current iteration $t$ of Algorithm 2. Since $\sigma_{\text{exact}}$ is unknown in real situations, a possible strategy for its estimation will be discussed in the next section. In Table 3, we report the values of TP, FN, FP, Jaccard index and RMSE of the reconstructed signals for both models. The final estimated $\lambda$ is also reported in Table 3, together with the number of performed homotopy iterations and of the value $\sigma_{\text{target}}$. Using homotopy, we observe that we retrieve values which are comparable with the best ones obtained using SFW with grid search. This shows the effectiveness of the homotopy strategy. Overall, the algorithm applied to solve ($\tilde{\mathcal{D}}_{KL}-|\cdot|$) yields better results than ($L^{2}-|\cdot|$) in the presence of Poisson noise, with a reduction of the number of FN and an improvement of the accuracy in terms of Jaccard index.
@@ 5.2 Simulated 2D and 3D examples: choice of $\sigma_{\text{target}}$
Table 4: Different estimates of $\sigma_{\text{target}}$ in the 2D simulated setting